Backlinks to DiaryRun12 in ARGOS Web (Search all webs)

Results from ARGOS web retrieved at 23:37 (GMT)

Commissioning
Commissioning run diaries Run 1st day Run ID Description 20111128 Run # 1 Nov 2011 Installation of dichroic interfaces 20130614 Run Installati...
DayTime20160510
13:30: Arrival at telescope 16:00: we try to start up both systems. Due to power outage the SX LGSW computer was off, we turned it on again. The DX LGSW computer ...
r4 - 11 May 2016 - 21:50 by MarcoBonaglia
DayTime20160511
LgswUnitSx and LgswUnitDx HW servicing 8:00: we unpack the PCD and the amplifier from 3L boxes, we bring it to 5. We also spend 1:30h in searching for the WfsCame...
DayTime20160512
5:30: Gustavo deploys the arm, we have pbl starting AdSecSx: it periodically faild while loading the DSP program. We try many times to power cicle. LB finally can...
r6 - 13 May 2016 - 15:16 by TommasoMazzoni
DayTime20160513
TT w/ Pyr Bin4 test 8:00: we start with PyrWfs, setup as Bin4, mag 15. * CCD47 is not set to Bin16 during the preset * TS at Bin4 is missing the KL basis at...
r4 - 24 May 2016 - 12:03 by MarcoBonaglia
DayTime20160514
Laser Power Test Measured sum power: * SX 43.256 * DX 48.631 SIDE L1 L2 L3 DX Field cam 15.50 18.06 18.17 DX Calorimeter 15.59 17.06 15.63 SX ...
r9 - 15 May 2016 - 07:33 by TommasoMazzoni
DayTime20160515
Fast guiding check w/ oscilloscope 11:30: Tommaso and Sebastian go on the telescope to check the Fast Guiding delay. We connect an oscilloscope after the 10x divi...
r8 - 24 May 2016 - 12:24 by MarcoBonaglia
NightTime20160515
18:00: back to SX, we setup for the night We start testing the new telescope offset procedure. We have a first try and the AOS hangs: Mar May 16 03:05:56 DX syste...
r2 - 28 Oct 2016 - 12:12 by MarcoBonaglia
NightTime20160516
20:00: we are on sky with lasers 20:30: FLAO fails, CCD39 was off at 3:29:49, we had an error in the FYSOM state machine for OPT.LOOPBUF NOMAD 1499_0210548 * ...
NEW - 24 May 2016 - 12:57 by MarcoBonaglia
NightTime20160517
DT: training of XY on SW and HW LgswUnitSx calibrations 20:30: we deploy the cal unit and we try to do SX calibrations SX HEXA 2.79 0.32 0.78 208 3...
NEW - 24 May 2016 - 13:45 by MarcoBonaglia
NightTime20160518
BCU SX debugging 15:00: we start to check the behaviour of the SX BCU. We notice that the main issue is when the Pyramid slopes are enabled. On DX we measure slop...
r3 - 20 May 2016 - 00:50 by MarcoBonaglia
NightTime20160519
20:00: problems w/ Sx AdSec, it goes in panic mode at each restart. SDSS J13432.85 415503 20:39: AdSec started. Propagate lasers. TN AO Gain AO CL TT Res...
r2 - 01 May 2017 - 06:19 by MarcoBonaglia
NightTime20160520
03:15UT: doing flats. SDSSJ1110 6459 03:45UT: loop closed. TN AO Gain AO CL TT Res. WF nm rms HO Res. WF nm rms AO rec DIMM " JCL PCL N...
NightTime20160521
02:46UT: getting ready, spotters out. SDSSJ1110 6459 03:00UT: telescope collimated. 03:09UT: loop closed, including spot optimization, at first try! usual mask a...
NightTime20160522
Late opening for technical reasons SDSSJ1110 6459 04:10UT Loop closed 04:24UT We keep on being visited by airplanes 04:45UT Un po di snapshot al passare delle n...
NightTime20160523
Laser head exchange DX Laser 2 with Head #4 was exchanged by Head #5 Laser power measurement. DX Laser 1 Laser 2 Laser 3 Field measurement 16.01 1...
NightTime20160524
03:00UT: Si sale a controllare il DichroicSX che non si muove correttamente. Non si capisce bene, sembra che il Fold Mirror duri fatica a partire. Nel pomeriggio ...
Number of topics: 17
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback